OpenStudy (anonymous):

Write the equation of the line that passes through (–3, 4) with a slope of 2 in point-slope form

OpenStudy (anonymous):

y-y1=m(x-x1) substitute the given

OpenStudy (anonymous):

Write the equation of the line that passes through (–3, 4) with a slope of 2 in point-slope form. a)y – 4 = 2(x + 3) b) y + 4 = 2(x – 3) c)y + 3 = 2(x – 4) d)y – 3 = 2(x + 4)

OpenStudy (anonymous):

y-4=2(x--3) y-4=2(x+3)

OpenStudy (anonymous):

a

OpenStudy (anonymous):

thank you